# Text-to-Model: Text-Conditioned Neural Network Diffusion for Train-Once-for-All Personalization

### v1

@@ 1 Introduction
Generative artificial intelligence (GenAI) has been flourishing in different aspects of human life, and people can simply generate content from natural language text prompts [1, 2, 3, 4]. Large language models [1, 5], like GPT-4, have especially shown emergent intelligence [6] in the knowledge of language through text-to-text transformation [7, 8, 1, 5]. Besides, recent progress in text-to-image (e.g., stable diffusion) [9, 4, 2, 10] and text-to-video (e.g., Sora) [3, 11] diffusion models has shown the great power of AI in understanding the physical world and generating high-quality images and videos that are virtually indistinguishable from reality [12, 3]. The text-prompted GenAI maps the human languages’ semantics to the world knowledge in different forms in language and vision. One step further, in this paper, we propose and study whether the GenAI can understand hyper-level knowledge—the knowledge inherently resides in the AI itself models’ parameters. Specifically, we study text-to-model generation; akin to text-to-text, text-to-image, and text-to-video, text-to-model targets whether the GenAI models can directly generate the model parameters given the human’s text prompts to meet the personalization demand of diverse end users.
@@
We focus on a practical scenario called train-once-for-all personalization [13], which means that the generic model is trained just once and can later be customized into a condensed model on the fly for different end-users and requests, given their task descriptions. For example, the CIFAR-100 dataset [14] contains 100 classes, but an end user may just need a personalized model with a certain 10 classes according to a specific scenario (e.g., classifying items in the kitchen). In other words, train-once-for-all personalization targets that train the model once and customize the model to be well performed in a sub-distribution when deployed, and an example is in Figure 1. But there are tremendous sub-distributions, for the CIFAR-100 example, the number of personalized 10-way tasks is (${100\choose 10}=1.73\times 10^{13}$, even not taking permutations into consideration, so it is challenging for the GenAI model to generalize. Inspired by recent progress in neural network diffusion [15, 16], we propose Tina, a Text-Conditioned Neural Network Diffusion for Train-Once-for-All Personalization. Tina is trained on model parameters with the models’ task descriptions, and it can be generalized to unseen tasks, or even unseen classes (entities), given the text prompts.
@@
In Tina, a CLIP model [17] is used to embed the users’ task descriptions into the diffusion model as the conditions. The diffusion model of Tina is the diffusion transformer (DiT) [12] that is shown to have high expressive power under scaling law in the fields of image [12] and video generation [3]. We demonstrate that DiT’s scaling law applies to model parameter generation as well: increasing the number of parameters and data sizes enhances the model’s capability to generalize across more challenging tasks that involve scaling the dimension of generated models. However, it is surprising to find that even though the number of personalized tasks is astronomical (e.g., $1.73\times 10^{13}$ for 10-way tasks), by our designs, Tina can generalize on extremely small datasets ($\sim 1000$ data points) and support different lengths of classification tasks (5-way or 8-way tasks, etc.) in training once. Our analysis shows that Tina can reach both in-distribution and out-of-distribution personalization of generated models. Thanks to the vision-language alignment of CLIP, Tina can also take images as prompts and generalize under few-shot or even zero-shot settings. We also verify whether Tina understands world knowledge by testing its abilities under prompts of natural language descriptions and predicting unseen entities. Our contributions are as follows:
@@ 2.1.1 Definition of Setup
Following [13], we consider image classification for train-once-for-all personalization due to the natural personalization requirements of image classification. We note that our method is not limited to classification tasks and can be extended to other tasks for personalization. Define a task $k$ as classification over a subset of classes $\mathcal{Y}_{k}\subset\mathcal{Y}$. The goal of personalization is to learn a neural network predictor $f_{\theta_{k}}:\mathcal{X}\mapsto\mathcal{Y}_{k}$, parameterized by $\theta_{k}$. To handle many tasks at the same time, we further assume we have the task description natural text $t_{k}$ for $\mathcal{Y}_{k}$, and it is generally the description of the classes and styles of $\mathcal{Y}_{k}$. We want to build a neural network generator $G(t_{k})$ where given $t_{k}$, it will output the model parameters $\theta_{k}$. Specifically, consider using a large-scale dataset with many classes covering $\mathcal{Y}$ to learn the personalized-friendly function $f_{\theta_{k}}=G_{\phi}(t_{k})$ parameterized by $\phi$. $G_{\phi}$ is learned on the large dataset to generate any personalized model directly from the task descriptions, and the setup is called train-once-for-all personalization [13]. Train-once-for-all personalization has wide applications in a server-user system, where the model generator $G_{\phi}$ is learned on the server for personalized cloud services to many future users. We refer to [13] for more detailed advantages and usages of train-once-for-all personalization.
@@ 2.1.2 Strong Baselines: Classifier Selection and TAPER
Classifier Selection. For a generic network $f_{\theta}$, we consider that it consists of a feature extractor parameterized by $\psi$ with a linear classifier $\mathbf{w}=[\mathbf{w}^{(1)},\ldots,\mathbf{w}^{(|\mathcal{Y}|)}]$ of $|\mathcal{Y}|$ vectors for output predictions over all classes in $\mathcal{Y}$. The generic model is trained on the large dataset, and we want to personalize it into a few-way classification task $k$. One effective method is to build a personalized classifier $\mathbf{w}_{k}$ by selecting only the row vectors in $\mathbf{w}$ for the relevant classes. Therefore, the personalized model for task $k$ are $\theta_{k}=\{\psi,\mathbf{w}_{k}\}$, and this approach is called classifier selection, which serves as a strong baseline [13].
TAPER. We briefly introduce TAPER [13] proposed by the original paper on train-once-for-all personalization and discuss its limitations. The main idea of TAPER is to train several experts (bases) and learn a mixture network to fuse these experts into a personalized model. It has three stages as follows.
@@ 2.1.3 Dataset Preparation and Description
Testing data preparation. The overall demonstration of data partitions can be found in Figure 2. The blue blocks refer to the training data, and the green blocks are the testing data. For testing, there are two kinds of evaluation metrics: i) In-distribution (ID, the light green blocks): the personalized tasks are seen during training of the generative model $G$, and $G$ generates the p-Models tested on the testset of each seen task. ii) Out-of-distribution (OOD, the dark green blocks): the tasks are unseen during the generator $G$’s training, and $G$ directly generates the p-Models from the task prompts (the text descriptions). We note that the original TAPER cannot be tested on the OOD tasks since it requires the target personalized training data for finetuning the expert models. To remedy this, we derive TAPER-Mixer to only train the mixer without finetuning the experts and verify its OOD generalization on unseen tasks.
@@ 2.2 Proposed Tina: Text-conditioned Neural Network Diffusion Model
Figure 3: Framework overview of Tina.
@@ 2.2.1 Framework Overview
2.2.2 Architecture and Training Objective
We use diffusion models as the generative model and follow the main architecture of G.pt [16] that uses a diffusion transformer as the backbone. Analogous to the optimization process that takes random initialization as inputs and outputs the trained models, the diffusion process takes the noise as inputs and gradually denoises to recover the original distributions. Previous works have shown the rationale of neural network diffusion [16, 15, 18]. We choose DiT as the backbone because it can be easily scaled up and is shown to have great generalization and expressiveness. We use signal prediction for the diffusion process and inherit the architecture of GPT-2 [8] as the transformer. The used text encoder is the pretrained ViT-B/32 in CLIP [17].
@@ 2.2.2 Architecture and Training Objective
where the timestep $j$ is embedded in DiT by frequency-based encoding [19]. The detailed training procedure is in Algorithm 1. We use DDPM sampling [9]; add Gaussian noise depicted by the $\bar{\alpha}$ to $\theta_{k}$ and gradually denoising it.
@@ 2.2.3 Design Details
Data augmentation. In [16], the permutation invariance property [20, 21, 22] is utilized for data augmentation by randomly permuting the neurons without changing the function. However, in our scenario, we find this augmentation will even impede training. We hypothesize that the personalized models are finetuned from the same generic model, so they may lie in the same or close loss landscape basins; as a result, permutation augmentation will disturb network representations and impair Tina training. Further, we develop an effective classifier augmentation strategy to speed up Tina training under limited data by randomly permuting the order of classes in the task description and also the order of corresponding classifier vectors during training. This data augmentation improves sample diversity and helps the DiT better learn the description-to-classifier sequence modeling in a position-aware manner.
Parameter inheritance. In [16], the authors release a pretrained checkpoint of G.pt, which is also DiT for parameter generation. G.pt is pretrained on large datasets of optimization checkpoints; though it has different conditions, designs, and scenarios from ours, we explore whether we can inherit some parameters from the pretrained checkpoints to speed up and boost training. Considering the model sizes and architectures are different, we use a strategy similar to bert2BERT [23, 24, 25] for inheriting parameters.
Classification sequence padding. We study how to incorporate more personalized settings where diverse users request for tasks with different numbers of classes. In language models [26, 5], padding is used to enable sequence-to-sequence learning with different input and output lengths. Inspired by this, we use the padding technique to enable the description-to-classifier sequence of different classification lengths. Specifically, if the user’s number of classes is smaller than the maximal length, we pad missing classes with tokens ‘<->’ in the task description list and mask the corresponding classifier vectors with zero-like tensors. We denote this strategy as classification sequence padding, and Tina can learn to adapt to any number of classes within the maximal length.
@@ 3.1 Experimental Setups
Datasets and p-Models. We use three datasets to conduct experiments: Mini-ImageNet [27, 28], CIFAR-100 [14], and Caltech-101 [29]. Mini-ImageNet is a subset of the ImageNet dataset, primarily used for few-shot learning tasks. CIFAR-100 is a popular benchmark dataset for image classification tasks. Each class contains 600 images, divided evenly into 20 superclasses and 100 classes. Caltech-101: A dataset for object recognition featuring diverse images with varied resolutions and quality. It includes 101 categories, each containing 40 to 800 images, offering a wide range of objects and scenes compared to CIFAR-100 and Mini-ImageNet. For the images with different resolutions, we resize them into 32 $\times$ 32 for unified modeling. The personalized tasks are crafted by selecting 10 classes out of the 100/101 total classes. If not mentioned otherwise, the number of p-Models (i.e., personalized tasks) for training Tina is 1000.
We use two architectures for personalized models: a simple CNN (dubbed as CNN) and ResNet-20 (dubbed as ResNet). The CNN architecture follows [16], which consists of 2 layers, and the number of parameters is approximately 5K. We take all the parameters of CNN as the input and output of Tina. But for ResNet-20, the number of parameters is nearly 272k, which is too large for Tina’s generation. Thus, we explore partial parameter generation following [15]. We only personalize the classifier layers for parameter generation, nearly 640 parameters.
@@
Compared baselines. We follow the baselines used in the original paper of train-once-for-all personalization [13]. As described in subsection 2.1.3, we use the generic model trained in stage 1 as a baseline, showing the performance without any personalization. Further, we compare the classifier selection method described in subsection 2.1.2, which serves as a strong baseline for personalization [13]. The vanilla TAPER [13] requires finetuning the expert models on the target tasks and cannot generalize on out-of-distribution personalization where only target text descriptions are available. For fair comparisons, we adopt TAPER-Mixer, which adopts the mixer of TAPER for generating the aggregation weights, and the MLP-based mixer can generalize on unseen tasks.
Evaluation metrics. For Table 1, we compare in-distribution personalization and out-of-distribution personalization as elaborated in subsection 2.1.3. For other tables and figures, we report the out-of-distribution personalization as p-Acc.
@@ 3.2 Results under Different Datasets
In Table 1, we evaluate the performance of our proposed method, Tina, against several baseline methods including Generic Model, Classifier Selection, and TAPER-Mixer across various datasets and model architectures for the task of train-once-for-all personalization. It is found that the Generic Model has inadequate performance, validating the need for personalization techniques. For the personalization methods, the results demonstrate that Tina consistently outperforms all baseline methods across both in-distribution and out-of-distribution personalization scenarios. Though Tina is a text-to-model foundation model, it is worth noting that Tina shows intelligence of personalization under limited data (nearly 1000 samples). Specifically, for in-distribution personalization, Tina achieves significant improvements with an average score of 79.94, surpassing the next best method, Classifier Selection, by a margin of 3.19. Similarly, for out-of-distribution personalization, Tina leads with an average score of 80.55, which is a notable increase over the second-best performing method by 2.78. It is notable that TAPER-Mixer shows performance gains over Classifier Selection in CNN but has marginal results in ResNet. Also, TAPER-Mixer has inferior performance compared with Tina, showing the advantages of Tina as a generative model in parameter generation. TAPER-Mixer only learns to merge the expert models, while Tina learns to directly generate the parameters.

### v2

@@ 1 Introduction
Generative artificial intelligence (GenAI) has been flourishing in different aspects of human life, and people can simply generate content from natural language text prompts (Brown et al., 2020; Ramesh et al., 2022; OpenAI, 2024; Rombach et al., 2022). Large language models (Brown et al., 2020; Touvron et al., 2023), like GPT-4, have especially shown emergent intelligence (Bubeck et al., 2023) in the knowledge of language through text-to-text transformation (Radford et al., ; Radford et al., 2019; Brown et al., 2020; Touvron et al., 2023). Besides, recent progress in text-to-image (e.g., stable diffusion) (Nichol & Dhariwal, 2021; Rombach et al., 2022; Ramesh et al., 2022; Zhang et al., 2023) and text-to-video (e.g., Sora) (OpenAI, 2024; Singer et al., 2023) diffusion models has shown the great power of AI in understanding the physical world and generating high-quality images and videos that are virtually indistinguishable from reality (Peebles & Xie, 2023; OpenAI, 2024). The text-prompted GenAI maps the human languages’ semantics to the world knowledge in different forms in language and vision. One step further, in this paper, we propose and study whether the GenAI can understand hyper-level knowledge—the knowledge inherently resides in the AI itself models’ parameters. Specifically, we study text-to-model generation; akin to text-to-text, text-to-image, and text-to-video, text-to-model targets whether the GenAI models can directly generate the model parameters given the human’s text prompts to meet the personalization demand of diverse end users.
@@
We focus on a practical scenario called train-once-for-all personalization (Chen et al., 2023), which means that the generic model is trained just once and can later be customized into a condensed model on the fly for different end-users and requests, given their task descriptions. For example, the CIFAR-100 dataset (Krizhevsky, 2009) contains 100 classes, but an end user may just need a personalized model with a certain 10 classes according to a specific scenario (e.g., classifying items in the kitchen). In other words, train-once-for-all personalization targets that train the model once and customize the model to be well performed in a sub-distribution when deployed, and an example is in Figure 1. But there are tremendous sub-distributions, for the CIFAR-100 example, the number of personalized 10-way tasks is (${100\choose 10}=1.73\times 10^{13}$, even not taking permutations into consideration, so it is challenging for the GenAI model to generalize. Inspired by recent progress in neural network diffusion (Wang et al., 2024; Peebles et al., 2022), we propose Tina, a Text-Conditioned Neural Network Diffusion for Train-Once-for-All Personalization. Tina is trained on model parameters with the models’ task descriptions, and it can be generalized to unseen tasks, or even unseen classes (entities), given the text prompts.
@@
In Tina, a CLIP model (Radford et al., 2021) is used to embed the users’ task descriptions into the diffusion model as the conditions. The diffusion model of Tina is the diffusion transformer (DiT) (Peebles & Xie, 2023) that is shown to have high expressive power under scaling law in the fields of image (Peebles & Xie, 2023) and video generation (OpenAI, 2024). We demonstrate that DiT’s scaling law applies to model parameter generation as well: increasing the number of parameters and data sizes enhances the model’s capability to generalize across more challenging tasks that involve scaling the dimension of generated models. However, it is surprising to find that even though the number of personalized tasks is astronomical (e.g., $1.73\times 10^{13}$ for 10-way tasks), by our designs, Tina can generalize on extremely small datasets ($\sim 1000$ data points) and support different lengths of classification tasks (5-way or 8-way tasks, etc.) in training once. Our analysis shows that Tina can reach both in-distribution and out-of-distribution personalization of generated models. Thanks to the vision-language alignment of CLIP, Tina can also take images as prompts and generalize under few-shot or even zero-shot settings. We also verify whether Tina understands world knowledge by testing its abilities under prompts of natural language descriptions and predicting unseen entities. Our contributions are as follows:
@@ 2.1.1 Definition of Setup
Following Chen et al. (2023), we consider image classification for train-once-for-all personalization due to the natural personalization requirements of image classification. We note that our method is not limited to classification tasks and can be extended to other tasks for personalization. Define a task $k$ as classification over a subset of classes $\mathcal{Y}_{k}\subset\mathcal{Y}$. The goal of personalization is to learn a neural network predictor $f_{\theta_{k}}:\mathcal{X}\mapsto\mathcal{Y}_{k}$, parameterized by $\theta_{k}$. To handle many tasks at the same time, we further assume we have the task description natural text $t_{k}$ for $\mathcal{Y}_{k}$, and it is generally the description of the classes and styles of $\mathcal{Y}_{k}$. We want to build a neural network generator $G(t_{k})$ where given $t_{k}$, it will output the model parameters $\theta_{k}$. Specifically, consider using a large-scale dataset with many classes covering $\mathcal{Y}$ to learn the personalized-friendly function $f_{\theta_{k}}=G_{\phi}(t_{k})$ parameterized by $\phi$. $G_{\phi}$ is learned on the large dataset to generate any personalized model directly from the task descriptions, and the setup is called train-once-for-all personalization (Chen et al., 2023). Train-once-for-all personalization has wide applications in a server-user system, where the model generator $G_{\phi}$ is learned on the server for personalized cloud services to many future users. We refer to Chen et al. (2023) for more detailed advantages and usages of train-once-for-all personalization.
@@ 2.1.2 Strong Baselines: Classifier Selection and TAPER
Classifier Selection. For a generic network $f_{\theta}$, we consider that it consists of a feature extractor parameterized by $\psi$ with a linear classifier $\mathbf{w}=[\mathbf{w}^{(1)},\ldots,\mathbf{w}^{(|\mathcal{Y}|)}]$ of $|\mathcal{Y}|$ vectors for output predictions over all classes in $\mathcal{Y}$. The generic model is trained on the large dataset, and we want to personalize it into a few-way classification task $k$. One effective method is to build a personalized classifier $\mathbf{w}_{k}$ by selecting only the row vectors in $\mathbf{w}$ for the relevant classes. Therefore, the personalized model for task $k$ are $\theta_{k}=\{\psi,\mathbf{w}_{k}\}$, and this approach is called classifier selection, which serves as a strong baseline (Chen et al., 2023).
TAPER. We briefly introduce TAPER (Chen et al., 2023) proposed by the original paper on train-once-for-all personalization and discuss its limitations. The main idea of TAPER is to train several experts (bases) and learn a mixture network to fuse these experts into a personalized model. It has three stages as follows.
@@ 2.1.3 Dataset Preparation and Description
Testing data preparation. The overall demonstration of data partitions can be found in Figure 2. The blue blocks refer to the training data, and the green blocks are the testing data. For testing, there are two kinds of evaluation metrics: i) In-distribution (ID, the light green blocks): the personalized tasks are seen during training of the generative model $G$, and $G$ generates the p-Models tested on the testset of each seen task. ii) Out-of-distribution (OOD, the dark green blocks): the tasks are unseen during the generator $G$’s training, and $G$ directly generates the p-Models from the task prompts (the text descriptions).
@@ 2.2 Proposed Tina: Text-conditioned Neural Network Diffusion Model
Figure 3: Framework overview of Tina. (a) Training stage. The p-Models are firstly augmented by our classifier augmentation strategy and then noised according to the diffusion step. The p-Models are tokenized into chunks of vectors, and the classification sequence padding is optionally used if the classification length is shorter than the default. The CLIP text encoder is used to encode the users’ text prompts during training. (b) Testing stage. Random noises are tokenized and denoised into parameters of p-Models. Thanks to the vision-language alignment of CLIP, Tina takes both text and visual prompts as the diffusion conditions.
@@ 2.2.1 Framework Overview
2.2.2 Architecture and Training Objective
We use diffusion models as the generative model and follow the main architecture of G.pt (Peebles et al., 2022) that uses a diffusion transformer as the backbone. Analogous to the optimization process that takes random initialization as inputs and outputs the trained models, the diffusion process takes the noise as inputs and gradually denoises to recover the original distributions. Previous works have shown the rationale of neural network diffusion (Peebles et al., 2022; Wang et al., 2024; Yuan et al., 2024). We choose DiT as the backbone because it can be easily scaled up and is shown to have great generalization and expressiveness. We use signal prediction for the diffusion process and inherit the architecture of GPT-2 (Radford et al., 2019) as the transformer. The used text encoder is the pretrained ViT-B/32 in CLIP (Radford et al., 2021).
@@ 2.2.2 Architecture and Training Objective
where the timestep $j$ is embedded in DiT by frequency-based encoding (Mildenhall et al., 2021). The detailed training procedure is in Algorithm 1. We use DDPM sampling (Nichol & Dhariwal, 2021); add Gaussian noise depicted by the $\bar{\alpha}$ to $\theta_{k}$ and gradually denoising it.
@@ 2.2.3 Design Details
Data augmentation. In Peebles et al. (2022), the permutation invariance property (Entezari et al., 2022; Ainsworth et al., 2023; Li et al., 2024b) is utilized for data augmentation by randomly permuting the neurons without changing the function. However, in our scenario, we find this augmentation will even impede training. We hypothesize that the personalized models are finetuned from the same generic model, so they may lie in the same or close loss landscape basins; as a result, permutation augmentation will disturb network representations and impair Tina training. Further, we develop an effective classifier augmentation strategy to speed up Tina training under limited data by randomly permuting the order of classes in the task description and also the order of corresponding classifier vectors during training. This data augmentation improves sample diversity and helps the DiT better learn the description-to-classifier sequence modeling in a position-aware manner.
Parameter inheritance. In Peebles et al. (2022), the authors release a pretrained checkpoint of G.pt, which is also DiT for parameter generation. G.pt is pretrained on large datasets of optimization checkpoints; though it has different conditions, designs, and scenarios from ours, we explore whether we can inherit some parameters from the pretrained checkpoints to speed up and boost training. Considering the model sizes and architectures are different, we use a strategy similar to bert2BERT (Chen et al., 2015; ; Qin et al., 2022) for inheriting parameters.
Classification sequence padding. We study how to incorporate more personalized settings where diverse users request for tasks with different numbers of classes. In language models (Devlin et al., 2018; Touvron et al., 2023), padding is used to enable sequence-to-sequence learning with different input and output lengths. Inspired by this, we use the padding technique to enable the description-to-classifier sequence of different classification lengths. Specifically, if the user’s number of classes is smaller than the maximal length, we pad missing classes with tokens ‘<->’ in the task description list and mask the corresponding classifier vectors with zero-like tensors. We denote this strategy as classification sequence padding, and Tina can learn to adapt to any number of classes within the maximal length.
@@ 3.1 Experimental Setups
Datasets and p-Models. We use three datasets to conduct experiments: Mini-ImageNet (Deng et al., 2009; Vinyals et al., 2016), CIFAR-100 (Krizhevsky, 2009), and Caltech-101 (Fei-Fei et al., 2004). Mini-ImageNet is a subset of the ImageNet dataset, primarily used for few-shot learning tasks. CIFAR-100 is a popular benchmark dataset for image classification tasks. Each class contains 600 images, divided evenly into 20 superclasses and 100 classes. Caltech-101: A dataset for object recognition featuring diverse images with varied resolutions and quality. It includes 101 categories, each containing 40 to 800 images, offering a wide range of objects and scenes compared to CIFAR-100 and Mini-ImageNet. For the images with different resolutions, we resize them into 32 $\times$ 32 for unified modeling. The personalized tasks are crafted by selecting 10 classes out of the 100/101 total classes. If not mentioned otherwise, the number of p-Models (i.e., personalized tasks) for training Tina is 1000.
We use two architectures for personalized models: a simple CNN (dubbed as CNN) and ResNet-20 (dubbed as ResNet). The CNN architecture follows Peebles et al. (2022), which consists of 2 layers, and the number of parameters is approximately 5K. We take all the parameters of CNN as the input and output of Tina. But for ResNet-20, the number of parameters is nearly 272k, which is too large for Tina’s generation. Thus, we explore partial parameter generation following Wang et al. (2024). We only personalize the classifier layers for parameter generation, nearly 640 parameters.
@@
Compared baselines. We follow the baselines used in the original paper of train-once-for-all personalization (Chen et al., 2023). As described in subsection 2.1.3, we use the generic model trained in stage 1 as a baseline, showing the performance without any personalization. Further, we compare the classifier selection method described in subsection 2.1.2, which serves as a strong baseline for personalization (Chen et al., 2023). We use TAPER as the important baseline.
Evaluation metrics. For Table 1, we compare in-distribution personalization and out-of-distribution personalization as elaborated in subsection 2.1.3. For other tables and figures, we report the out-of-distribution personalization as p-Acc. It is notable that for every setting, we test the personalization performances across over 100 tasks (i.e., 100 independent trials) and take the average scores presented in the tables and figures, and each task includes more than 1000 testing image samples. Therefore, the evaluation measurement is statistical, representative, and fair for the compared methods.
@@ 3.2 Results under Different Datasets
In Table 1, we evaluate the performance of our proposed method, Tina, against several baseline methods including Generic Model, Classifier Selection, and TAPER across various datasets and model architectures for the task of train-once-for-all personalization. It is found that the Generic Model has inadequate performance, validating the need for personalization techniques. For the personalization methods, the results demonstrate that Tina consistently outperforms all baseline methods across both in-distribution and out-of-distribution personalization scenarios. Though Tina is a text-to-model foundation model, it is worth noting that Tina shows intelligence of personalization under limited data (nearly 1000 samples). Specifically, for in-distribution personalization, Tina achieves significant improvements with an average score of 79.94, surpassing the next best method, Classifier Selection, by a margin of 3.19. Similarly, for out-of-distribution personalization, Tina leads with an average score of 80.55, which is a notable increase over the second-best performing method by 2.78. It is notable that TAPER shows performance gains over Classifier Selection in CNN but has marginal results in ResNet. Also, TAPER has inferior performance compared with Tina, showing the advantages of Tina as a generative model in parameter generation. TAPER only learns to merge the expert models, while Tina learns to directly generate the parameters.
Table 2: Results on larger p-Models (ViT-B/32).
Dataset CIFAR-100 Caltech-101
In-distribution Personalization
Classifier Selection 95.07 96.33
TAPER 95.25 96.32
Tina 95.45 97.15
Out-of-distribution Personalization
Classifier Selection 94.78 96.08
TAPER 94.96 96.15
Tina 95.15 96.72
Results on larger and more complex p-Models. We verify whether Tina is effective for larger and more complex p-Models in Table 2. We use ViT-B/32 pretrained by CLIP and Tina generates the personalized layers in ViT as described for ResNet. The results are promising that our Tina can reach 97.15’s accuracy in personalization when using the SOTA backbone ViT-B/32 pretrained by CLIP, and Tina also consistently outperforms the baselines. It showcases the scalability and potential of Tina to be adopted in trending and SOTA architectures and reach SOTA performances in personalization.